In 1955, Life Magazine reported that a 25-year-old mother of three worked, on average, an 80 hour week. Recently, many groupsE Part (c) In words, state what your random variable X represents X represents the number of hours a woman works in a week X0 Part (f What is the p-value? (Round your answer to four decimal places.) Explain what the p-value means for this problem OIndicate the correct decision (reject or do not reject the null hypothesis), the reason for it, and write an appropriate cPart (i) Construct a 95% confidence interval or he rue mean Sketch the graph o the situation. Label upper bounds to two decimPART F PART i
Answer #1

The hypotheses are

Ha : μ>80

Given the sample size n=68 and sample mean X 83 , sample standard deviation s=10

The test statistic is

X-80 83- 80 10 V68 t- 2.4738634

f) The P-value of the test is

P-value P(T> 2.4738634) P-value 0.0080

When \alpha =0.05 ,

Reject the null hypothesis since P-value 0.0080 a 0.05

The conclusion is "There is sufficient evidence to conclude that the average number of hours women work each week is more than 80 hours".

i) The -a) 10090 two sided confidence interval for mean based on the sample data is

\overline{x}\pm t_{1-\alpha /2,n-1}\frac{s}{\sqrt{n}}

Here the 95% confidence interval is

10 83 ± t1-0.05/2,67 V68 (80.58,85.42)